# BEYOND CLASSICAL AND CONTEMPORARY MODELS: A TRANSFORMATIVE AI FRAMEWORK FOR STUDENT DROPOUT PREDICTION IN DISTANCE LEARNING USING RAG, PROMPT ENGINEERING, AND CROSS-MODAL FUSION

### v1

@@ 1. Introduction
To capture temporal regularities in raw click streams, researchers progressed from static tabular learners to deep sequence architectures. Early bidirectional LSTMs and GRUs improved recall on week by week logs, but Temporal Convolutional Networks (TCNs) soon emerged as a faster alternative with comparable accuracy, thanks to causal dilations that expand the receptive field while preserving order [8]. The current benchmark, I TCN MOOC, adds channel wise attention and reaches an AUC of 0.90 on 9 153 learners six points above XGBoost on the same split [16] . Transformer variants such as ED Transformer exploit self attention to model semester long dependencies and report state of the art macro F1 of 0.84 on a 40 000 student open university corpus [17] . In parallel, graph neural networks (GNNs) recast the platform as a heterogeneous graph whose nodes represent students, resources, and forum threads. Models such as GAT4Edu propagate information along ”co view” or ”reply to” edges, capturing peer influence and content affinity. These approaches outperform sequence nets when interaction density is high (e.g., AUC = 0.92 on a 60 000 node edX forum graph) but struggle when cohorts number fewer than 5 000, where the graph is too sparse to learn robust embeddings. Both sequence and graph models remain largely black boxes, seldom incorporate affective text, and rarely deliver instance level explanations gaps that motivate the interpretable, multimodal design of TRIAD Drop. Discussion forum posts, chat transcripts, and reflective essays form a rich affective layer that complements behavioral logs. Early studies counted emotion lexicon terms ’confused,’ ’bored,’ and so on to show that negative tone predicts withdrawal. Transformer models now deliver far finer resolution. Yang, Wen, and Rosé fine tuned BERT on 1.2 million Coursera comments and found that anxiety laden language raises dropout odds by 22 % (odds ratio = 1.22, $p<0.01$) [18]. Sun et al. then modeled sentiment trajectories with RoBERTa, demonstrating that a shift from curiosity to frustration two weeks before the census date lifts the hazard ratio to 1.36 on edX data [19], Alario-Hoyos et al. observed comparable patterns in Spanish vocational MOOCs with the multilingual XLM-R model, underscoring the cross-lingual generality of affective signals [20]. Despite these advances, textual cues are usually bolted on post hoc: most studies concatenate sentiment scores or topic proportions to tabular vectors and feed them to XGBoost instead of learning a joint representation with click streams. Temporal alignment how emotion and engagement co-evolve and instance-level attribution which sentence triggered the alert therefore remain under-explored. Closing those gaps requires architectures that fuse linguistic and behavioral modalities end-to-end and can point to the specific utterances driving a risk prediction, capabilities targeted in TRIAD-Drop via gated cross-modal attention and RAG-grounded explanations. Standard LLMs excel at free-text generation but often hallucinate. Lewis et al. mitigated this with RAG-Sequence, which lets the decoder attend to passages fetched from a dense-vector index, cutting factual errors in open-domain QA by 45 % [10]. Industrial variants followed: IBM’s Fabric-RAG adds adaptive re-ranking and reports a 57 % hallucination drop in financial-report summarization. Education is beginning to adopt the idea on the conversational side. Nguyen and Hale grafted RAG onto an academic-advising chatbot, boosting student trust scores by 0.6 Likert points and halving unanswered queries [21] . Zhou et al. recently released ’RAG-Tutor,” which grounds formative feedback in course handbooks and lifts perceived helpfulness to 92 % in a 600-learner study, while Shi and Sun showed curriculum-aligned answer generation with a prerequisite knowledge graph[22, 23] . Yet all current educational RAG systems are descriptive: they answer questions or give feedback after students act. None integrates retrieved evidence directly into predictive dropout models, nor do they couple those predictions with source-cited rationales. Bridging that gap by fusing RAG explanations with multimodal behavioral predictors is precisely the aim of TRIAD-Drop. Cross-attention transformers have reshaped multimodal research, models such as CLIP align images and captions via joint contrastive learning and routinely beat late-fusion baselines in computer vision and clinical triage [12] . Education is following suit. Sun et al. showed that simply concatenating RoBERTa-based sentiment scores with VLE logs and feeding the vector to an MLP lifts early-dropout recall by five percentage points, but offers no insight into which channel drives the decision [19] . Moving toward true fusion, EduFusion introduces gated cross-modal attention over four streams demographics, clicks, forum text, and video engagement achieving macro-F1 = 0.85 on a 40 000-student Chinese open-university data set, yet still lacks explanatory output [24] . Li et al. combine behavior graphs with comment embeddings via a hierarchical transformer and reach AUC = 0.91 on the public MOOCCube benchmark, but require heavy hyper-tuning and supply only global feature importances [25]. These studies confirm that learning adaptive weights across channels beats naïve stacking, but none pairs the fused predictor with instance-level, source-cited rationales. TRIAD-Drop advances the state of the art by integrating gated cross-attention for modality weighting and a RAG module that grounds every alert in retrieved policy or feedback passages, unifying predictive accuracy with actionable inter-pretability.
@@ 3.2. Dataset
Augmentation Methodology: To enable multimodal analysis, we synthetically generated student comment data using GPT-4.5 (gpt-4-turbo-2024-04-09 snapshot) through a structured prompting approach:
  1. (1)
    Feature-guided generation: For each student, we created 5 context-specific prompts incorporating their academic status, demographics, and performance metrics
        "As a [age]-year-old [gender] [marital status] student with
        [scholarship status] studying [course], who scored [grade]
        in [course unit], write a forum post about [random academic topic]"
  2. (2)
    Academic realism: Prompts incorporated domain-specific scenarios (assignment struggles, scheduling conflicts) validated by 3 subject matter experts
  3. (3)
    Diversity enforcement: Generated comments spanned emotional tones (frustration, confusion, satisfaction) proportional to academic performance
  4. (4)
    Quality control: 10% of generated comments were manually verified against authenticity metrics (PAPA scale [29])
The resulting augmented dataset contains 22,115 synthetic comments (5 per student) stored in the student_comment field. The complete augmented dataset and generation scripts are available in our project repository: https://github.com/Milouden/TRIAD-Drop.git.

### v2

@@ 1. Introduction
To capture temporal regularities in raw click streams, researchers progressed from static tabular learners to deep sequence architectures. Early bidirectional LSTMs and GRUs improved recall on week by week logs, but Temporal Convolutional Networks (TCNs) soon emerged as a faster alternative with comparable accuracy, thanks to causal dilations that expand the receptive field while preserving order [8]. The current benchmark, I TCN MOOC, adds channel wise attention and reaches an AUC of 0.90 on 9 153 learners six points above XGBoost on the same split [16] . Transformer variants such as ED Transformer exploit self attention to model semester long dependencies and report state of the art macro F1 of 0.84 on a 40 000 student open university corpus [17] . In parallel, graph neural networks (GNNs) recast the platform as a heterogeneous graph whose nodes represent students, resources, and forum threads. Models such as GAT4Edu propagate information along ”co view” or ”reply to” edges, capturing peer influence and content affinity. These approaches outperform sequence nets when interaction density is high (e.g., AUC = 0.92 on a 60 000 node edX forum graph) but struggle when cohorts number fewer than 5 000, where the graph is too sparse to learn robust embeddings. Both sequence and graph models remain largely black boxes, seldom incorporate affective text, and rarely deliver instance level explanations gaps that motivate the interpretable, multimodal design of TRIAD Drop. Discussion forum posts, chat transcripts, and reflective essays form a rich affective layer that complements behavioral logs. Early studies counted emotion lexicon terms ’confused,’ ’bored,’ and so on to show that negative tone predicts withdrawal. Transformer models now deliver far finer resolution. Yang, Wen, and Rosé fine tuned BERT on 1.2 million Coursera comments and found that anxiety laden language raises dropout odds by 22 % (odds ratio = 1.22, $p<0.01$) [18]. Sun et al. then modeled sentiment trajectories with RoBERTa, demonstrating that a shift from curiosity to frustration two weeks before the census date lifts the hazard ratio to 1.36 on edX data [19], Alario-Hoyos et al. observed comparable patterns in Spanish vocational MOOCs with the multilingual XLM-R model, underscoring the cross-lingual generality of affective signals [20]. Despite these advances, textual cues are usually bolted on post hoc: most studies concatenate sentiment scores or topic proportions to tabular vectors and feed them to XGBoost instead of learning a joint representation with click streams. Temporal alignment how emotion and engagement co-evolve and instance-level attribution which sentence triggered the alert therefore remain under-explored. Closing those gaps requires architectures that fuse linguistic and behavioral modalities end-to-end and can point to the specific utterances driving a risk prediction, capabilities targeted in TRIAD-Drop via gated cross-modal attention and RAG-grounded explanations. Standard LLMs excel at free-text generation but often hallucinate. Lewis et al. mitigated this with RAG-Sequence, which lets the decoder attend to passages fetched from a dense-vector index, cutting factual errors in open-domain QA by 45 % [10]. Industrial variants followed: IBM’s Fabric-RAG adds adaptive re-ranking and reports a 57 % hallucination drop in financial-report summarization. Education is beginning to adopt the idea on the conversational side. Nguyen and Hale grafted RAG onto an academic-advising chatbot, boosting student trust scores by 0.6 Likert points and halving unanswered queries [21] . Zhou et al. recently released ’RAG-Tutor,” which grounds formative feedback in course handbooks and lifts perceived helpfulness to 92 % in a 600-learner study, while Shi and Sun showed curriculum-aligned answer generation with a prerequisite knowledge graph[22, shi2023curriculum] . Yet all current educational RAG systems are descriptive: they answer questions or give feedback after students act. None integrates retrieved evidence directly into predictive dropout models, nor do they couple those predictions with source-cited rationales. Bridging that gap by fusing RAG explanations with multimodal behavioral predictors is precisely the aim of TRIAD-Drop. Cross-attention transformers have reshaped multimodal research, models such as CLIP align images and captions via joint contrastive learning and routinely beat late-fusion baselines in computer vision and clinical triage [12] . Education is following suit. Sun et al. showed that simply concatenating RoBERTa-based sentiment scores with VLE logs and feeding the vector to an MLP lifts early-dropout recall by five percentage points, but offers no insight into which channel drives the decision [19] . Moving toward true fusion, EduFusion introduces gated cross-modal attention over four streams demographics, clicks, forum text, and video engagement achieving macro-F1 = 0.85 on a 40 000-student Chinese open-university data set, yet still lacks explanatory output [24] . Li et al. combine behavior graphs with comment embeddings via a hierarchical transformer and reach AUC = 0.91 on the public MOOCCube benchmark, but require heavy hyper-tuning and supply only global feature importances [25]. These studies confirm that learning adaptive weights across channels beats naïve stacking, but none pairs the fused predictor with instance-level, source-cited rationales. TRIAD-Drop advances the state of the art by integrating gated cross-attention for modality weighting and a RAG module that grounds every alert in retrieved policy or feedback passages, unifying predictive accuracy with actionable inter-pretability.
@@ 3.2. Dataset
The complete dataset and generation scripts are available in our project repository: https://github.com/Milouden/TRIAD-Drop.git.